OpenStudy (anonymous):

Are 12/40 and 9/30 equivalent or not equivalent.

OpenStudy (anonymous):

12/40=3/10 9/30= 3/10 they are equivalent

OpenStudy (anonymous):

yes both equal to 3/10

OpenStudy (anonymous):

You can cross multiply too and see if you get the same answer

OpenStudy (anonymous):

oh ok. i get it now

OpenStudy (anonymous):